A pure titanium cube has an edge length of 2.71in How many titanium atoms does it contain? Titanium has a density of 4.50 g/cm^3.
Answer #1

2.7 in =6.858 cm
volume = 6.858^3 =322.547 cm3
mass =ρxV =4.5x322.547 =1451.46 gm

atomic mass of Ti= 47.87 gm

no of moles of Ti=1451.46/47.87 = 30.32

no. of atoms of Ti = no. of moles x avogadro no. = 30.32x(avogadro no.) = 1.826E25
Answer #2

edge length = 2.7*2.54 =6.858 cm
volume =322.546 cm^3
mass of cube =1451.459 g
no of atoms =1451.459/47.867 = 30.3
no of atoms ≈ 30.3